# On second order conditions for singular optimal control of port-Hamiltonian systems

### v2

@@ Abstract
We study singular optimal control problems of port-Hamiltonian (descriptor) systems. We employ general control-affine cost functionals that include as a special case the energy supplied to the system. We first derive optimality conditions for the case of ordinary differential equations with and without control bounds by applying the general theory to the specially structured port-Hamiltonian case, and show that this leads to nice optimality conditions, in particular in the linear case. We then extend these results to classes of port-Hamiltonian descriptor systems.
Keywords: singular optimal control, port-Hamiltonian descriptor system, Goh condition, Legendre-Clebsch condition, minimal supplied energy.
AMS subject classification: 49J15, 37J99, 34A09
@@ 1 Introduction
The energy based modeling of physical systems employing the model class of port-Hamiltonian descriptors, see e.g. [29, 43], has become an accepted modeling paradigm which has been successfully used in a multitude of applications from a wide variety of application domains: mechanics [5, 6, 24, 42], electrical engineering [37, 41], thermodynamics and fluid dynamics [2, 13, 35, 36], economics [23]. See [9, 14, 43] and the recent survey [29].
In this paper we consider a simplified form of real nonlinear port-Hamiltonian (pH) descriptor systems, see [28, 29], that have a governing equation
@@
with the state $x$ belonging to the state space $\mathcal{X}:=W^{1,1}([0,T];\mathbb{R}^{n})$ (the Sobolev space of functions in $L^{1}[0,T]$ with weak first derivatives that have a finite $L^{1}$-norm) and the input (control) $u$ ranging in the control space $\mathcal{U}:=L^{1}([0,T];\mathbb{R}^{m})$ with $L^{1}$ denoting the usual Lebesgue space. Equation (1.1) has an associated energy function
$\mathcal{E}:\mathbb{R}^{n}\to\mathbb{R},\quad\ x\mapsto\frac{1}{2}x^{T}E^{T}Qx$ (1.2)
(often called Hamiltonian or storage function) that satisfies
$\frac{\partial\mathcal{E}}{\partial x}(x)=E^{T}Qx.$ (1.3)
The output $y$ is given by
$y=G^{T}(x)Qx.$ (1.4)
As is common in pH (descriptor) systems, the structure of the coefficients is the following:
@@
  • $E,Q\in\mathbb{R}^{n,n}$ are such that $E^{T}Q$ is symmetric and positive semidefinite and represents the Hessian of $\mathcal{E}=\frac{1}{2}x^{T}E^{T}Qx$;
@@
Note that if $E$ is invertible, then one obtains the classical form of a pH system by a change of variables (see [43]).
The success of modeling within the class of pH (descriptor) systems is due to its many important properties which include the invariance of the class under power-conserving interconnection, which greatly simplifies modularized automated modeling, and the invariance under Galerkin projection which makes the class very suitable for discretization and model reduction. Physical properties like energy dissipation, Lyapunov stability and passivity are encoded in the structure of the equations leading directly to another key property of port-Hamiltonian (descriptor) systems, the power balance equation and the resulting dissipation inequality, see e.g. [28, 29] or [43] for the case of ordinary pH systems. From a physics point of view, pH systems model the interaction of three types of energy/power. The stored energy (in the energy storing components) is presented by the nonnegative quadratic form $\mathcal{E}(x)$, the dissipated energy by the nonnegative quadratic form $\mathcal{D}(x)=-x^{T}Q^{T}RQx$ and the supplied energy by the form $\mathcal{S}(y,u)=y^{T}u$.
@@
A classical example of pH modeling are (dissipative) Hamiltonian equations of motion, see e.g. [1, 43], which, in first order descriptor representation using the position coordinate $q$ and velocity coordinate $\tilde{p}$ (replacing the usual linear momentum $p=M\dot{q}$), together with a force term $\tilde{B}(q)u$, take the descriptor form
@@
i.e., kinetic plus potential energy, where $M$ is a positive (semi-)definite mass matrix, $K$ is a positive definite stiffness matrix, and $D(q,p)=D(q,p)^{T}$ is a positive semidefinite coefficient matrix that models internal damping. This representation has the advantage that small masses can be set to 00, leading to a singular mass matrix $M$. See [30] for a detailed analysis of different representations and transformations between the different formulations. A typical optimal control application is arising in the control of the system into an equilibrium state, e.g. in the stabilization of buildings excited by earthquakes, see e.g. [25].
With an increased understanding of the many advantageous properties of modeling with pH (descriptor) systems, in recent years also feedback control and optimal control problems for this model class have become an important research topic in different application fields [8, 10, 11, 17, 19, 26, 31, 32, 38, 39, 45].
@@
[…]
It is readily available as a mathematical expression in terms of the ports and it is interesting from an application point of view to employ it as the objective to be minimized in optimal control problems (OCPs) [10, 11, 32, 39]. However, using the supplied energy as cost function, the resulting OCP is typically singular, it is missing the (commonly used) positive quadratic regularization term in $u$, and standard solution techniques, for instance the construction of a Riccati state feedback, see e.g. [21, 27], are not directly applicable. In view of this difficulty, the topic of this paper is the analysis of singular OCPs with constraints given by pH (descriptor) systems without the regularization approach.
We obtain our results by applying optimality conditions for singular optimal control of control-affine systems. More precisely, considering problems of the form
@@
with $f_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, for $i=0\ldots,m$ and $\Psi:\mathbb{R}^{n}\to\mathbb{R},$ one gets the well-known Goh conditions which state that, along optimal trajectories,
$p^{T}[f_{i},f_{j}]=0,\quad\text{for }i,j=1,\dots,m,$ (1.10)
where $p$ is the adjoint state, and where, for any pair of smooth vector fields $h,k\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$, $[h,k]:=k^{\prime}h-h^{\prime}k$ is their Lie bracket, with $k^{\prime},h^{\prime}$ denoting the Jacobians of $k,h$. Moreover, the generalized Legendre-Clebsch conditions ensure that the matrix with entries
$p^{T}[f_{j},[f_{0},f_{i}]]$ (1.11)
is negative semidefinite and, whenever it is definite, it provides a feedback formula for the singular control. These results are detailed in Appendix A.
@@
At this point we recall that any problem with a general control-affine integral cost can be rewritten in the Mayer form (1.7)-(1.9) by adding an extra scalar state variable (see (2.1) below). Therefore, the formulation (1.7)-(1.9) covers problems with integral-type costs of the form $\mathcal{J}\coloneqq\int_{0}^{T}\ell_{0}(x)+\sum_{i=1}^{m}\ell_{i}(x)u_{i}\,dt,$ where $\ell_{i}:\mathbb{R}^{n}\to\mathbb{R}$ for $i=0\ldots,m$.
The paper is organized as follows.
In Appendix A we recall second order optimality conditions for problems of the form (1.7)-(1.9). In Section 2 we analyze how these general optimality conditions look when we specialize the constraint function from general ordinary differential equations to ordinary port-Hamiltonian systems and we show that the structure helps to obtain much more elegant conditions. We extend these results to port-Hamiltonian descriptor systems in Section 3 and we present some conclusions and directions of future work in Section 4.
2 Optimality conditions for nonlinear pH descriptor systems
@@
The optimality conditions for singular optimal control problems that are presented in Appendix A do not exploit the pH structure. In this section, we therefore make explicit use of the nonlinear pH structure in (1.1) and generalize results of [10, 26, 39].
@@ 2.1 Optimality conditions for ordinary nonlinear pH systems
In this case, recalling the Hamilton function introduced in (A.1), the adjoint variable has an additional coordinate $p_{n+1}$ and takes the form $\begin{bmatrix}p\\ p_{n+1}\end{bmatrix},$ where $p$ maps into $\mathbb{R}^{n}$ and $p_{n+1}$ is scalar. In the absence of (state or terminal) constraints, the multiplier $p_{n+1}$ is unique and the associated extremal is normal, i.e., the corresponding multiplier $p_{n+1}$ can be set to $1$. Therefore, throughout the remainder of the article, we fix $p_{n+1}=1$.
Theorem 2.1 (Goh conditions for nonlinear ordinary pH control).
@@
$\left\{\begin{array}[]{rl}p^{T}[g_{i},g_{j}]&=0,\\ \ell_{j}^{\prime}\,g_{i}-\ell_{i}^{\prime}\,g_{j}&=0,\end{array}\right.\qquad\text{for }\,i,j=1,\ldots,m,$
@@
The result follows from Theorem A.1 Appendix A. ∎
Using latter result, we compute the expressions for the generalized Legendre-Clebsch conditions (see Theorem A.2 in Appendix A, and $W$ and $d$ given in (A.8)) in the context of problem (1.7)–(1.9). We have
$[f_{0},f_{i}]=\begin{bmatrix}[g_{0},g_{i}]\\ \ell_{i}^{\prime}g_{0}-\ell_{0}^{\prime}g_{i}\end{bmatrix},$ (2.4)
@@
$[f_{j},[f_{0},f_{i}]]=\\ \begin{bmatrix}[g_{j},[g_{0},g_{i}]]\\ \ell_{i}^{\prime\prime}g_{0}g_{j}+\ell_{i}^{\prime}g_{0}^{\prime}g_{j}-\ell_{0}^{\prime\prime}g_{i}g_{j}-\ell_{0}^{\prime}g_{i}^{\prime}g_{j}-\ell_{j}^{\prime}[g_{0},g_{j}]\end{bmatrix}.$
@@
$w_{ij}\coloneqq p^{T}[g_{j},[g_{0},g_{i}]]\\ +\Big{(}\ell_{i}^{\prime\prime}g_{0}g_{j}+\ell_{i}^{\prime}g_{0}^{\prime}g_{j}-\ell_{0}^{\prime\prime}g_{i}g_{j}-\ell_{0}^{\prime}g_{i}^{\prime}g_{j}-\ell_{j}^{\prime}[g_{0},g_{i}]\Big{)},$ (2.5)
and
$d_{i}\coloneqq p^{T}[g_{0},[g_{0},g_{i}]]\\ +\Big{(}\ell_{i}^{\prime\prime}g_{0}g_{0}+\ell_{i}^{\prime}g_{0}^{\prime}g_{0}-\ell_{0}^{\prime\prime}g_{i}g_{0}-\ell_{0}^{\prime}g_{i}^{\prime}g_{0}-\ell_{0}^{\prime}[g_{0},g_{i}]\Big{)},$ (2.6)
@@
Theorem 2.2 (Legendre-Clebsch condition for nonlinear ordinary pH control).
For the nonlinear pH optimal control problem (2.1), the matrix $W=[w_{ij}],$ with $w_{ij}$ given in (2.5), is negative semidefinite. Additionally, whenever $W$ is negative definite over some control interval then, on that interval, one can write the optimal control in feedback form via the expression
$Wu+d=0.$ (2.7)
@@
The proof follows from Theorem A.2 in Appendix A. ∎

### v3

@@ Abstract
We study nonlinear singular optimal control problems of port-Hamil-tonian (descriptor) systems. We employ general control-affine cost functionals that include as a special case the energy supplied to the system. We first derive optimality conditions for the case of ordinary differential equations with and without control bounds by applying the general theory to the specially structured port-Hamiltonian case, and show that this leads to elegant optimality conditions, in particular in the linear case. We then extend these results to classes of nonlinear port-Hamiltonian descriptor systems.
@@ 1 Introduction
The energy-based modeling of physical systems employing the model class of port-Hamiltonian (descriptor) systems, see e.g. [MehU23, SchaJelt2014], has become an accepted paradigm which has been successfully used in a multitude of applications from a wide variety of practical domains: mechanics [brugnoli2019por1t, brugnoli2019port2, macchelli2007port, siuka2011port], electrical engineering [reis2021analysis, schoberl2008modelling], thermodynamics and fluid dynamics [Altmann2017, hoang2011port, rashad2021port, rashad2021port2], economics [macchelli2014towards]. See [duindam2009modeling, Jacob2012, SchaJelt2014] and the recent survey [MehU23].
In this paper we consider a simplified form of real nonlinear port-Hamiltonian (pH) descriptor systems, see [MehM19_CDC, MehU23], that have a governing equation
@@
with an associated energy function
$\mathcal{E}:\mathbb{R}^{n}\to\mathbb{R},\quad\ x\mapsto\frac{1}{2}x^{T}E^{T}Qx,$ (1.2)
(often called Hamiltonian or storage function). As is common in pH (descriptor) systems, the structure of the coefficients is the following:
@@
  • $E,Q\in\mathbb{R}^{n,n}$ are such that $E^{T}Q$ is symmetric and positive semidefinite and represents the Hessian of $\mathcal{E}(x)=\frac{1}{2}x^{T}E^{T}Qx$, i.e., $\frac{\partial\mathcal{E}}{\partial x}(x)=E^{T}Qx$;
@@
For (1.1)-(1.2), the output $y$ is given by
$y=G^{T}(x)Qx.$ (1.3)
Note that if $E$ is invertible, then one obtains the classical form of a pH system by a change of variables (see [SchaJelt2014]).
The success of modeling within the class of pH (descriptor) systems is due to its many important properties which include the invariance of the class under power-conserving interconnection, which greatly simplifies modularized automated modeling, and the invariance under Galerkin projection which makes the class very suitable for discretization and model reduction, see [CarHLMR24, RasCSS20]. Physical properties like energy dissipation, Lyapunov stability and passivity are encoded in the structure of the equations leading directly to another key property of port-Hamiltonian (descriptor) systems, the power balance equation and the resulting dissipation inequality, see e.g. [MehM19_CDC, MehU23] or [SchaJelt2014] for the case of ordinary pH systems. From a physics point of view, pH systems model the interaction of three types of energy/power. The stored energy (in the energy storing components) is represented by the nonnegative quadratic form $\mathcal{E}(x)$, the dissipated energy by the nonnegative quadratic form $\mathcal{D}(x)=-x^{T}Q^{T}RQx$ and the supplied energy by the form $\mathcal{S}(y,u)=y^{T}u$.
@@
A classical example of pH modeling are (dissipative) Hamiltonian equations of motion, see e.g. [AbrM08, SchaJelt2014], which, in first order descriptor representation using the position coordinate $q$ and velocity coordinate $\tilde{p}$ (replacing the usual linear momentum $p=M\dot{q}$), together with a force term $\tilde{B}(q)u$, take the descriptor form
@@
i.e., kinetic plus potential energy, where $M$ is a positive (semi-)definite mass matrix, $K$ is a positive semidefinite stiffness matrix, and $D(q,p)=D(q,p)^{T}$ is a positive semidefinite coefficient matrix that models internal damping. This representation has the advantage that small masses can be set to 0, leading to a singular mass matrix $M$. See [MehS23] for a detailed analysis of different representations and transformations between the different formulations. A typical optimal control application is arising in the control of the system into an equilibrium state, e.g. in the stabilization of buildings excited by earthquakes, see e.g. [MahZS00, SchZBSTW24].
With an increased understanding of the many advantageous properties of modeling with pH (descriptor) systems, in recent years also feedback control and optimal control problems for this model class have become an important research topic. While the theoretical analysis from different perspectives is studied in [FauKMPSW25, FauMPSW22, wu2020reduced], and we will come back to this later, the use in different application fields is studied, e.g., in dynamic network flow problems [DogKLST23], learning control [KoeSSH21, NagLJB15], electrothermal microgrids [KriS21], and thermodynamics [Masc22].
@@
$\mathcal{J}(x,u)=\int_{t_{0}}^{T}\mathcal{S}(y(\tau),u(\tau))\,d\tau=\int_{t_{0}}^{T}x(\tau)^{T}Q^{T}G(x)u(\tau)\,d\tau.$
It is readily available as a mathematical expression in terms of the ports and it is interesting from an application point of view to employ it as the objective to be minimized in optimal control problems (OCPs) [FauKMPSW25, FauMPSW22, lhmnc], where as cost function one typically uses
$\tilde{J}(x,u):=\frac{1}{2}\begin{bmatrix}x^{T}&u^{T}\end{bmatrix}\begin{bmatrix}\tilde{Q}&\tilde{S}\\ \tilde{S}^{T}&\tilde{R}\end{bmatrix}\begin{bmatrix}x\\ u\end{bmatrix}$.
with semidefinite $\begin{bmatrix}\tilde{Q}&\tilde{S}\\ \tilde{S}^{T}&\tilde{R}\end{bmatrix}$. While it is in general no problem to drop the quadratic term in $x$, i.e. by setting $\tilde{Q}=0$ [Meh91], it is usually not possible to set $\tilde{R}=0$, as is done when the supplied energy is used as cost function. In this case the resulting OCP is typically singular, since it is missing the (commonly used) positive quadratic regularization term in $u$, and standard solution techniques, for instance the construction of a Riccati state feedback, see e.g. [KunM11a, KunM24, Meh91], are not directly applicable. In view of this difficulty, the topic of this paper is the analysis of singular OCPs with constraints given by pH (descriptor) systems without the regularization approach, i.e. without the term $u^{T}\tilde{R}u$ in the cost function.
The paper is organized as follows. In Section 2 we recall second order optimality conditions for general control-affine problems. In Section 3, we analyze how these optimality conditions look when we specialize the constraint function from general control-affine ordinary differential equations to ordinary port-Hamiltonian systems. We show that the structure helps to obtain much more elegant conditions. We extend these results to port-Hamiltonian descriptor systems in Section 4 and we present some conclusions and directions of future work in Section 5.
2 On Goh and generalized Legendre-Clebsch conditions
We obtain our results by applying optimality conditions for singular optimal control of control-affine systems, known as Goh and generalized Legendre-Clebsch conditions, that we briefly introduce in the sequel. We consider problems of the form
@@ 2 On Goh and generalized Legendre-Clebsch conditions
with $f_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, for $i=0\ldots,m$ and $\Psi:\mathbb{R}^{n}\to\mathbb{R}.$ The state $x$ belongs to the state space $\mathcal{X}:=W^{1,\infty}([0,T];\mathbb{R}^{n})$ (the Sobolev space of functions in $L^{\infty}([0,T];\mathbb{R}^{m}),$ with weak first order derivatives that have a finite $L^{\infty}$-norm, with $L^{\infty}$ denoting the usual Lebesgue space) and the input (control) $u=(u_{1},\dots,u_{m})^{T}$ ranges in the control space $\mathcal{U}:=L^{\infty}([0,T];\mathbb{R}^{m})$. At this point we recall that any problem with a general control-affine integral cost can be rewritten in the Mayer form (2.1)-(2.3) by adding an extra scalar state variable (see (3.1) below). Therefore, the formulation (2.1)-(2.3) covers problems with integral-type costs of the form $\mathcal{J}\coloneqq\int_{0}^{T}\ell_{0}(x)+\sum_{i=1}^{m}\ell_{i}(x)u_{i}\,dt,$ where $\ell_{i}:\mathbb{R}^{n}\to\mathbb{R}$ for $i=0\ldots,m$.
For nonlinear systems as the one in (2.2), one cannot obtain the control straightforwardly from Pontryagin’s Maximum Principle (PMP). More precisely, the maximum condition provided by the PMP results in a maximization with respect to the control variable of the Hamilton function, which is itself a control-affine function
$\mathcal{H}=\lambda^{T}f=\lambda^{T}(f_{0}(x)+\sum_{i=1}^{m}f_{i}(x)u_{i}),$ (2.4)
$\lambda$ here being the adjoint state, whose associated adjoint equation is given by
$\dot{\lambda}^{T}=\partial_{x}\mathcal{H}=\lambda^{T}\partial_{x}f,\quad\lambda(T)=\nabla\Psi(x(T)).$ (2.5)
Consequently, whenever
$\mathcal{H}_{u_{i}}=\lambda^{T}f_{i}=0,$ (2.6)
holds along a non-trivial interval of time for some $i\in\{1,\dots,m\},$ no information about the optimal control can be retrieved from (2.4). When this situation occurs along an optimal trajectory, it is said that the corresponding control has a singular arc. Briefly speaking, singular arcs of a control component $u_{i}$ correspond to intervals in which $\frac{\partial\mathcal{H}}{\partial u_{i}}$ vanishes and the corresponding Hessian matrix $\frac{\partial^{2}\mathcal{H}}{\partial u^{2}}$ is singular (which holds trivially in the current control-affine framework). In the absence of control constraints, (2.6) results in a necessary condition for optimality. Since (2.6) does not depend explicitly on the control variable, one can differentiate with respect to time and obtains
$0=\frac{d}{dt}\mathcal{H}_{u_{i}}=\dot{\lambda}^{T}f_{i}+\lambda^{\top}f_{i}^{\prime}\dot{x}=\lambda^{T}\big([f_{0},f_{i}]+\sum_{j=1}^{m}u_{j}[f_{j},f_{i}]\big),$ (2.7)
where we used the state equation (2.2) and the adjoint equation (2.5) to replace $\dot{x}$ and $\dot{\lambda},$ and where $f_{i}^{\prime}$ denotes the Jacobian, i.e., the derivative of $f_{i}$ with respect to $x$, and for any pair of smooth vector fields $\colon\mathbb{R}^{n}\to\mathbb{R}^{n}, [h,k]:=kh$$k^{\prime}h-h^{\prime}k$ is their Lie bracket, with $k^{\prime},h^{\prime}$ denoting the Jacobians of $k,h$. A remarkable property, discovered by Goh [Goh66] (and further explored by Robbins [Rob67]), is that the coefficients of the control in the first time derivative (2.7) of $\mathcal{H}_{u_{i}}$ vanish. More precisely, Goh conditions state the following, see e.g. [ABDL12, FraTon13, schattler2012geometric] for rigorous proofs.
Theorem 2.1 (Goh condition).
For the optimal control problem (2.1)-(2.3), the Goh condition
$\lambda^{T}[f_{i},f_{j}]=0,\quad\text{for }i,j=1,\dots,m$ (2.8)
holds along optimal trajectories.
Remark 2.1.
At this point it is worth recalling that the traditional technique of adding a regularizing term of the form $\alpha\sum_{i=1}^{m}u_{i}^{2}$ with positive $\alpha$ and then taking $\alpha\to 0$ does not work in this context, since (2.6) will depend on $u$ and then the expression (2.7) would not be possible. On the other hand, if one aims at obtaining a feedback formula for the controls on singular arcs by applying Pontryagin’s Maximum Principle to the modified quadratic cost and taking $\alpha\to 0,$ the expression contains $\alpha$ in the denominator and the limit may fail to exist, see, e.g. [JacGL70]. Despite this difficulty, it has been shown in the case of linear constant coefficients that this problem has some hidden regularity in a subspace [FauKMPSW25].
The Goh property (2.8) guarantees then that (2.7) does not explicitly contain the control variables, which allows one to differentiate the expression (2.7) once more. In the second order derivative $\frac{d^{2}}{dt^{2}}\mathcal{H}_{u},$ the matrix coefficient of the control variable turns out to be semidefinite. Indeed, one gets
$\frac{d^{2}}{dt^{2}}\mathcal{H}_{u_{i}}=\lambda^{T}[f_{0},[f_{0},f_{i}]]+\sum_{j=1}^{m}u_{j}\lambda^{T}[f_{j},[f_{0},f_{i}]].$ (2.9)
Thus, from $\frac{d^{2}}{dt^{2}}\mathcal{H}_{u_{i}}=0$ one obtains a linear system for $u$ given by
$Wu+d=0,$ (2.10)
with the Hessian matrix $W=[w_{ij}]$ and the right hand side vector $d$, where for $i,j=1,\dots,m$,
$w_{ij}\coloneqq\lambda^{T}[f_{j},[f_{0},f_{i}]],\qquad d_{i}\coloneqq\lambda^{T}[f_{0},[f_{0},f_{i}]].$ (2.11)
This leads to a generalized Legendre-Clebsch condition, see e.g. [Krener1977, Theorem 6.2 and/or Corollary 6.3], [ABDL12], or [FraTon13].
Theorem 2.2 (Generalized Legendre-Clebsch condition).
For the nonlinear control-affine problem (2.1)-(2.3), the associated Hessian matrix $W$ in (2.11) is negative semidefinite along optimal trajectories. Additionally, if $W$ is actually negative definite, then, from (2.10), one can express the optimal control in feedback form.
Remark 2.2.
Note that, whenever $W$ is actually negative definite, the feedback obtained from (2.10) will depend on the state and adjoint variables.
The extensions of Theorems 2.1 and 2.2 to control-constrained problems are provided in the subsequent sections, in which we use Goh and the generalized Legendre-Clebsch conditions to deduce properties for pH systems.
3 Optimality conditions for nonlinear pH systems
@@
The optimality conditions for singular optimal control problems in the last section do not exploit the pH structure. In this section, we therefore make explicit use of the nonlinear pH structure in (1.1) and generalize results of [FauKMPSW25, Masc22, SchPFWM23].
@@ 3.1 Optimality conditions for ordinary nonlinear pH systems
In this case, recalling the Hamilton function introduced in (2.4), the adjoint variable has an additional coordinate $\lambda_{n+1}$ and takes the form $\begin{bmatrix}\lambda\\ \lambda_{n+1}\end{bmatrix}$, where $\lambda$ maps into $\mathbb{R}^{n}$ and $\lambda_{n+1}$ is scalar. In the absence of (state or terminal) constraints, the multiplier $\lambda_{n+1}$ is unique and the associated extremal is normal, i.e., the corresponding multiplier $\lambda_{n+1}$ can be set to $1$. Therefore, throughout the remainder of the article, we fix $\lambda_{n+1}=1$.
Remark 3.1.
Note that the use of the supplied energy as cost function in (3.1) corresponds to the special case in which $\ell_{0}(x)=0$ and $\ell_{i}(x)=y_{i}(x)$, for $i=1,\dots,m$. We will come back to this choice in more detail in the following section.
Corollary 3.1 (Goh conditions for nonlinear ordinary pH control).
@@
$\left\{\begin{array}[]{rl}\lambda^{T}[g_{i},g_{j}]&=0,\\ \ell_{j}^{\prime}\,g_{i}-\ell_{i}^{\prime}\,g_{j}&=0,\end{array}\right.\qquad\text{for }\,i,j=1,\ldots,m,$ (3.4)
@@
The result follows from Theorem 2.1. ∎
Using Corollary 3.4, we compute the expressions for the generalized Legendre-Clebsch conditions (see Theorem 2.2, and $W$ and $d$ given in (2.11)) in the context of problem (2.1)–(2.3). We have
$f_{0},f_{i}]=\begin{bmatrix}[g_{0},g_{i}]\\ \ell_{i}^{\prime}g_{0}-\ell_{0}^{\prime}g_{i}\end{bmatrix}$, (3.5)
@@
$f_{j},[f_{0},f_{i}]]=\begin{bmatrix}[g_{j},[g_{0},g_{i}]]\\ \ell_{i}^{\prime\prime}g_{0}g_{j}+\ell_{i}^{\prime}g_{0}^{\prime}g_{j}-\ell_{0}^{\prime\prime}g_{i}g_{j}-\ell_{0}^{\prime}g_{i}^{\prime}g_{j}-\ell_{j}^{\prime}[g_{0},g_{j}]\end{bmatrix}$.
@@
$w_{ij}\coloneqq\lambda^{T}[g_{j},[g_{0},g_{i}]]+\Big(\ell_{i}^{\prime\prime}g_{0}g_{j}+\ell_{i}^{\prime}g_{0}^{\prime}g_{j}-\ell_{0}^{\prime\prime}g_{i}g_{j}-\ell_{0}^{\prime}g_{i}^{\prime}g_{j}-\ell_{j}^{\prime}[g_{0},g_{i}]\Big),$ (3.6)
$d_{i}\coloneqq\lambda^{T}[g_{0},[g_{0},g_{i}]]+\Big(\ell_{i}^{\prime\prime}g_{0}g_{0}+\ell_{i}^{\prime}g_{0}^{\prime}g_{0}-\ell_{0}^{\prime\prime}g_{i}g_{0}-\ell_{0}^{\prime}g_{i}^{\prime}g_{0}-\ell_{0}^{\prime}[g_{0},g_{i}]\Big),$ (3.7)
@@
Corollary 3.2 (Generalized Legendre-Clebsch condition for nonlinear ordinary pH control).
For the nonlinear pH optimal control problem (3.1), the matrix $W=[w_{ij}],$ with $w_{ij}$ given in (3.6), is negative semidefinite. Additionally, whenever $W$ is negative definite over some control interval then, on that interval, one can write the optimal control in feedback form from (2.10).
@@
The proof follows from Theorem 2.2. ∎
Remark 3.2.
Note that when $W$ is non-singular, the control $u$ can be expressed in feedback form as a function of the state and costate variables. This representation is useful for establishing the regularity of the control and facilitates the application of numerical techniques, such as the shooting method.
While the Goh and generalized Legendre-Clebsch conditions from Corollaries 3.1 and 3.2 do not offer much insight for general nonlinear pH systems, they simplify significantly, or even hold trivially, in specific examples where symmetry properties naturally arise. This simplification reveals an interesting novelty, which is exemplified in the general nonlinear model for dissipative Hamiltonian equations of motion, introduced above in Example 1.1.
Example 3.3.
$\int_{0}^{T}y^{\top}u\,dt=\int_{0}^{T}\frac{\partial\mathcal{E}}{\partial p}^{\top}B(q)u\,dt=\int_{0}^{T}p^{\top}M^{-1}B(q)u\,dt$
the Lie brackets $f_{i},f_{j}]$ are given by
$f_{i},f_{j}]=\begin{bmatrix}0\\ 0\\ B_{j}(q)^{\top}\frac{\partial^{2}\mathcal{E}}{\partial p^{2}}B_{i}(q)-B_{i}(q)^{\top}\frac{\partial^{2}\mathcal{E}}{\partial p^{2}}B_{j}(q)\end{bmatrix}$.
These vanish identically for any sufficiently regular $\mathcal{E}$. In the current framework, one has $\frac{\partial^{2}\mathcal{E}}{\partial p^{2}}=M^{-1},$ for $M$ being symmetric positive definite, so Goh conditions (3.4) hold trivially. One gets the following expression for the second-order Lie brackets
$f_{j},[f_{0},f_{i}]]=\begin{bmatrix}B_{i}^{\prime}M^{-1}B_{j}+B_{j}^{\prime}M^{-1}B_{i}\\ 0\end{bmatrix}\\ +\begin{bmatrix}\frac{\partial D}{\partial p}(M^{-1}B_{i})B_{j}+\frac{\partial^{2}D}{\partial p^{2}}(M^{-1}p,B_{i})B_{j}\\ 0\end{bmatrix}$,